A boy throws a ball,B from top of tower down at velocity 20 m/s and another boy from the ground level of tower throws a ball,A up at 5m/sec.Height of tower is 100m.Find the time when balls will collide?
Why time when B will meet A is t=distance/velocity.?

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Relative velocity of both the balls=20m/s+5m/s=25m/s
Relative acceleration=0m/s^2
Initail distance between them 100m.
Hence time taken to collide=\dfrac{100m}{25m/s}=4s